# Human-In-The-Loop Machine Learning for Safe and Ethical Autonomous Vehicles: Principles, Challenges, and Opportunities

### v1

@@ V-C Key Findings and Insights
Ensuring accountability is also key in addressing the challenges. Accountability, in the context of AVs, is the framework of responsibilities, processes, and mechanisms that ensure these vehicles’ safe, ethical, and lawful operation. This includes clearly defining the roles of those involved, ensuring transparency and explainability of the AV’s decisions, monitoring and evaluating performance, taking corrective action when necessary, and enforcing consequences for non-compliance or failure. This framework aims to build trust, ensure safety, and maintain ethical standards in the introduction and operation of AVs.
Humans play a crucial role in ensuring accountability in AVs by providing responsive monitoring and performance evaluation. For example, in HITL-RL, humans can oversee the operation of RL algorithms, assess their performance, and implement corrective actions by injecting safe actions and preventing dangerous ones. If the RL system deviates from its target, humans can modify the reward model to guide and manage its operation better. XAI further enhances accountability by making the decision-making processes of AVs transparent and understandable. In AL, human annotators provide informative labels essential for AVs’ safe operation. By ensuring that the training data is accurately labeled, human annotators contribute to the robustness and reliability of AV systems. In CL, humans arrange tasks in order of increasing difficulty, providing safer and more manageable inputs for AVs. This method ensures a gradual learning process, progressively allowing the AV to build competence and safety.
@@
We observe HITL-ML can improve self-recognition and scene recognition in AVs to assist in managing complex situations and conducting ethical decisions in real-time. Moreover, humans play a crucial role in ensuring accountability in AVs by providing responsive monitoring and performance evaluation. Overall, integrating ethical considerations into the decision-making processes of AVs is critical to aligning these systems with societal norms and values. HITL-ML methods enhance the ethical decision-making capabilities of AVs by providing real-time adaptability, accountability, and improved detection capabilities. Incorporating HITL is essential for monitoring, evaluating and controlling AV operations to ensure ethical standards and safe, reliable performance in dynamic environments.
VI Future Trends and Research Directions
We investigate HITL-ML paradigms such as CL, HITL-RL, and AL, in conjunction with AVs. We also consider ethical AVs using HITL-ML. These approaches contribute to better learning ability and performance, safer and ethical systems, and more responsiveness and context aware in critical situations. These approaches enhance learning capabilities, improve performance, and create safer, more responsive systems in critical situations. However, there remain several opportunities to advance HITL-ML with AVs further.
@@ VI-B Timely Human Intervention
VI-C Security
Adverse conditions or potential cyber-attacks should not degrade the performance of HITL-ML for AVs, compromising their security and reliability. Security performance includes safeguarding human operators from malicious tampering and ensuring they can operate safely without causing unintended harm. Additionally, there is a risk of session hijacking, where an attacker takes over a session between a vehicle and a human operator. To prevent session hijacking, strong security measures such as encryption and two-way authentication are recommended. Authentication protocols should align with minimum overhead requirements, keep efficiency in message transmission, and support the unbounded number of message transmission issues in unreliable settings which is the case in AVs. Moreover, key agreement based on post-quantum cryptography and in particular Ring Learning with Errors (RLWE) is suggested to provide robust authentication for both humans and AVs. Overall, we need to establish rigorous testing, validation, and implementation of security protocols to prevent failures or exploits.
VI-D Regulatory Frameworks
The regulations for HITL-ML in AVs are designed to meet the public’s expectations around safety, legal responsibility, and privacy. The regulations ensure that HITL-ML systems in AVs are rigorously tested and meet high safety standards. How can comprehensive regulatory frameworks be developed and implemented to ensure the safe and effective integration of HITL processes into AV deployment, including defining the responsibilities for human operators, the safety standards and protocols for their intervention, and the ethical and legal boundaries of HITL systems? Moreover, how can we balance human effort and machine autonomy and regulate communication between humans and AVs to ensure effective status reporting and support in unforeseen situations? What are the most effective methods for documenting and evaluating the actions of both ML systems and human operators in HITL-ML frameworks for AVs to ensure appropriate legal accountability? How can regulations be designed to effectively protect human privacy in HITL-ML systems for AVs, ensuring transparent, consensual data collection and usage while preventing unauthorized access and breaches?
VI-E Safety
@@
Safety for HITL-ML in AVs includes a set of safety protocols, procedures and guidelines for human and AV operations. These safety measures are designed to ensure that HITL-ML in AVs are safe for users, passengers and pedestrians. At its core, the safety of HITL-ML in AVs is about using human commands, sensory data and other technologies to guide an AV through the streets while avoiding obstacles or potential traffic hazards. How can advanced safety systems be designed and implemented to combine human judgment and machine precision effectively? What specific roles should human operators play during critical situations? How can machine precision be used for real-time data processing, accurate decision-making, and consistent execution of safety protocols? How can these measures enhance the overall reliability and safety of AVs?
@@ VI-F Human Input
How can we ensure that human input is reliable, ethical, and consistent in these systems? How can the integration of human feedback into AI models of AVs be managed meticulously to avoid the introduction of human biases, ensuring that the subjective nature of human judgment does not skew the AI models away from objective decision-making? What constant vigilance and sophisticated validation mechanisms are needed to address these biases and ensure that human contributions enhance rather than detract from the model’s performance?
@@ VII Conclusion
We review the literature on HITL-ML for AVs, specifically we review the literature on CL and note how it helps organize the tasks in AVs from simple to complex and facilitates the learning process. Further, We review the literature on HITL-RL and find that its main benefit for AVs is to promote public trust and enhance safety. Moreover, we review the literature on AL and find that integrating AL with ML leads to better results in tasks such as anomaly detection, semantic mapping, object detection, and vehicle detection, contributing to the security of AVs. Moreover, we investigate the ethical implications of HITL-ML and the ways enhance self-recognition and scene recognition in AVs, assisting in managing complex situations. Finally, we specify the ways to further advance HITL-ML in AVs. Developing explainable AVs, timely human intervention, and regulatory frameworks along with tightenning security are critical for effective HITL-ML adoption.

### v2

@@ V-B Key Findings and Insights
However, another aspect emerges from the previous discussion, which is that of accountability and its importance in addressing the referred challenges. Accountability, in the context of AVs, is the framework of responsibilities, processes, and mechanisms that ensure these vehicles safe, ethical, and lawful operation. This includes clearly defining the roles of those involved, ensuring transparency and explainability of the AVs decisions, monitoring and evaluating performance, taking corrective action when necessary, and enforcing consequences for non-compliance or failure. This framework aims to build trust, ensure safety, and maintain ethical standards in the introduction and operation of AVs. Humans play a crucial role in ensuring accountability in AVs by providing responsive monitoring and performance evaluation.
@@
Overall, integrating ethical considerations into the decision-making processes of AVs is critical to aligning these systems with societal norms and values. HITL-ML methods enhance the ethical decision-making capabilities of AVs by providing real-time adaptability, accountability, and improved detection capabilities. Incorporating HITL is essential for monitoring, evaluating and controlling AV operations to ensure ethical standards and safe, reliable performance in dynamic environments.
VI Open Research Directions
In the previous sections we already pointed to several open issues emerging from the application of HITL-ML technologies to AVs. These technologies enhance learning capabilities, improve performance, and enable safer and more responsive systems in critical situations. In this section we bring up remaining opportunities to further advance HITL-ML globally within the scope of AVs.
@@ VI-B Timely Human Intervention
VI-C Safety
@@
Safety for HITL-ML in AVs includes a set of protocols, procedures and guidelines for humans and AV operations. These safety measures are designed to ensure that HITL-ML in AVs do not cause harm to users, passengers and pedestrians. At its core, the safety of HITL-ML in AVs is about using human commands, sensory data and other technologies to guide an AV through the streets while avoiding obstacles or potential traffic hazards. How can safe AV systems be designed and implemented combining human judgment and machine precision effectively? What specific roles should human operators play during critical situations? How can machine precision be used for real-time data processing, accurate decision-making, and consistent execution of safety protocols? How can these measures enhance the overall reliability and safety of AVs?
VI-D Security
Adverse conditions or potential cyber-attacks should not degrade the performance of HITL-ML for AVs, compromising their security and reliability. Security performance includes safeguarding human operators from malicious tampering and ensuring they can operate safely without causing unintended harm. Additionally, there is a risk of session hijacking, where an attacker takes over a session between a vehicle and a human operator. To prevent session hijacking, strong security measures such as encryption and two-way authentication are recommended. Authentication protocols should align with minimum overhead requirements, keep efficiency in message transmission, and cope with open unreliable media which is the case in AVs. Moreover, key agreement based on post-quantum cryptography and in particular Ring Learning with Errors (RLWE) is suggested to provide robust authentication for both humans and AVs. Overall, we need to establish rigorous testing, validation, and implementation of security protocols to prevent failures or exploits.
VI-E Regulatory Frameworks
The regulations for HITL-ML in AVs are designed to meet the public expectations around safety, legal responsibility, and privacy. The regulations ensure that HITL-ML systems in AVs are rigorously tested and meet high safety standards. How can comprehensive regulatory frameworks be developed and implemented to ensure the safe and effective integration of HITL processes into AV deployment, including defining the responsibilities for human operators, the safety standards and protocols for their intervention, and the ethical and legal boundaries of HITL systems? Moreover, how can we balance human effort and machine autonomy and regulate communication between humans and AVs to ensure effective status reporting and support in unforeseen situations? What are the most effective methods for documenting and evaluating the actions of both ML systems and human operators in HITL-ML frameworks for AVs to ensure appropriate legal accountability? How can regulations be designed to effectively protect human privacy in HITL-ML systems for AVs, ensuring transparent, consensual data collection and usage while preventing unauthorized access and breaches?
@@ VI-F Human Input
How can we ensure that human input is reliable, ethical, and consistent in these systems? How can the integration of human feedback into AI models of AVs be managed meticulously to avoid the introduction of human biases, ensuring that the subjective nature of human judgment does not skew the AI models away from objective decision-making? What constant vigilance and sophisticated validation mechanisms are needed to address these biases and ensure that human contributions enhance rather than detract model performance?
@@ VII Conclusion
In this paper we reviewed the literature on HITL-ML for AVs, specifically on CL, HITL-RL and AL technologies. We noted that CL helps organize the tasks in AVs from simple to complex and facilitates the learning process. HITL-RL promotes public trust and enhances safety. AL leads to better results in tasks such as anomaly detection, semantic mapping, object detection, and vehicle detection, also contributing to the safety of AVs. We then investigated the ethical implications of HITL-ML and surveyed the related literature concerning enhancing self-recognition and scene recognition in AVs to assist in managing complex situations. We also observed the importance of accountability in AVs operation. In all these topics we point to state-of-the-art works available in the literature. Finally, we go beyond the previous studies and enumerate a few global topics in HITL-ML applied to AVs that constitute current research opportunities. Namely, we believe that developing explainable AVs, enabling timely human intervention, enhancing safety and security, defining regulatory frameworks along with proper human input are critical for effective HITL-ML adoption.